OpenStudy (anonymous):

Serena bought 3/4 pound of cheese. She used 2/3 of the cheese for sandwiches. How much cheese did Serena use for sandwiches? A.1/4 lb B. 1/3 lb C. 1/2 lb D. 5/7 lb

OpenStudy (malcolmmcswain):

To find 2/3 of 3/4, we multiply: \[\frac{ 2 }{ 3 }*\frac{ 3 }{ 4 }=\frac{ 6 }{ 12 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

that is very helpful thank you sir :)

OpenStudy (malcolmmcswain):

No problem!